OpenStudy (anonymous):

Bayside insurance offers two health plans. Under plan A Giselle would have to pay the first $120 of her medical bills, plus 25% of the rest. Under plan B, Giselle would pay the first $210.00m but only 20% of the rest. For what amount of medical bills will plan B save Giselle money? Assume she has over $210.00 in bills. Giselle would save with plan B if she had more than $? in

OpenStudy (anonymous):

Plant A=0.25x+120 Plane B=0.2x+ 210 0.25x+120 = 0.2x+ 210 0.25x-0.2x= 210 -120 \[x \ge $1.800\]x= 90/0.05 x= S 1.800

OpenStudy (anonymous):

1,800?

OpenStudy (anonymous):

0.25x-0.2x= 210 -120 0.05x=90 x =90/0.05 x=1.800 yes